4. Investment Analysis Consider the evaluation of an agricultural investment A reqiring an initial outlay of S200,000 and will be evaluated over a 5-year period. It is assumed that there is no income tax obligations, no inflation, and no risk. Projected cash flows for this investment A are reported in the Table bolow. Time period 0 reflects the current date and the negativo $200,000 is the initial investment, and the rate of return is estimated at 8%. Year 0-present Investment A Co version Factor for (1.08) for-0.08 200,000 20,000 40,000 60,000 80,000 100,000 0.926 0.857 0.794 0.735 0.681 a) Using the inform ation above, calculate NPV of investment A knowing that b) Is this investment acceptable? Why or why not? Explain.
Answer #1

(a) Initial Investment = $ 200000

Total Present Value of Yearly Cash Flows = 20000 x 0.926 + 40000 x 0.857 + 60000 x 0.794 + 80000 x 0.735 + 100000 x 0.681 = $ 227340

NPV = 227340 - 200000 = $ 27340

(b) THe investment is acceptable as it generates a positive NPV for the firm undertaking the project.